satellite73 (satellite73):

if \(\sqrt{a}=c\sqrt{d}\) and \(\sqrt{b}=d\sqrt{c}\) then find \(\sqrt{a^2b^4}\) in terms of \(c\) and \(d\)

OpenStudy (anonymous):

jeez c'mon

OpenStudy (anonymous):

know i know what it is like to ask a question here c'mon what do you get?

OpenStudy (anonymous):

would it help to say that i have my answer, and i would like it checked?

OpenStudy (abb0t):

Please show your work. It is against CoC to give answers.

OpenStudy (anonymous):

lol

OpenStudy (abb0t):

:)

OpenStudy (anonymous):

problem is, i don't get the answers that poor @Big_Booty has in her (his?) choices

OpenStudy (anonymous):

thought i would get a second opinion it is not rocket science (brain surgery, whatever)

OpenStudy (anonymous):

you know, I dare not answer, I don't know what's going to pop up if I do

OpenStudy (kinggeorge):

Well \[\sqrt{a^2b^4}=ab^2\]\[a=c^2d\]\[b=d^2c\implies b^2=d^4c^2\]\[ab^2=c^4d^5\]is what I get.

OpenStudy (anonymous):

bless you

OpenStudy (anonymous):

I agree.

OpenStudy (anonymous):

that is what i got as well

OpenStudy (anonymous):

so now i don't feel like i screwed up and told @Big_Booty wrong no correct answer among the choices

OpenStudy (kinggeorge):

Seems like that problem was written terribly.

OpenStudy (anonymous):

yes, it does you want to see a great answer?

OpenStudy (anonymous):

It hapens =/ Its always the worst when you are helping someone and there is an error in their problem/text book. Then the person your helping starts to doubt if you are right, an argument can start, and you could start to doubt yourself o.O

OpenStudy (anonymous):

I got c4d5 as well

OpenStudy (anonymous):

someone asked how to write \(\frac{3}{\sqrt{8}}\) in simplest radical form, which they actually called "exact" form i said multiply top and bottom by \(\sqrt{2}\) and the teacher wrote back and said ESSAY FEEDBACK good start here but you don't have the square root of 16 you have square root of 64 and that equals 8. Then on the top you also need to multiply by square root of 8. The rule is you can not ever have a fraction with a square root denominator or zero or negative.

OpenStudy (anonymous):

still pissing me off i think i will go get some watermelon and take a break

OpenStudy (kinggeorge):

Reminds me of WebWork for calc 3. It was a good day when all the questions were written correctly.

OpenStudy (anonymous):

one thing for over worked underpaid grad students to screw up an routine question, far worse (imho) entirely for a "math teacher" to write that nonsense above

OpenStudy (anonymous):

most of ppl would just multiply square root of 8 tho

OpenStudy (anonymous):

and later divide by the common multiples and get the same answer u got hahaha

OpenStudy (anonymous):

look at this good start here but you don't have the square root of 16 you have square root of 64 really? did you look at the answer???

OpenStudy (anonymous):

thats plain stupidity lol

OpenStudy (anonymous):

and then says Then on the top you also need to multiply by square root of 8. like you have to believe it to see it!

OpenStudy (anonymous):

satellite's teacher

OpenStudy (anonymous):

and the final insult The RULE is you can not ever have a fraction with a square root denominator

OpenStudy (anonymous):

that being the answer to why \(\frac{3\sqrt{8}}{8}\) is the EXACT form of \(\frac{3}{\sqrt{8}}\)

OpenStudy (anonymous):

ok i will go take my xanax now and relax, maybe chase it with a shot of tulley

OpenStudy (anonymous):

i thought u were gonna get watermelon lol

OpenStudy (nincompoop):

a radical in the denominator is not "elegant" according to math gods and goddesses